# Physics-Inspired Synthesized Underwater Image Dataset

### v2

@@ Abstract
This paper introduces the physics-inspired synthesized underwater image dataset (PHISWID), a dataset tailored for enhancing underwater image processing through physics-inspired image synthesis. For underwater image enhancement, data-driven approaches (e.g., deep neural networks) typically demand extensive datasets, yet acquiring paired clean and degraded underwater images poses significant challenges. Existing datasets have limited contributions to image enhancement due to lack of physics models, publicity, and ground-truth images. PHISWID addresses these issues by offering a set of paired ground-truth (atmospheric) and underwater images synthetically degraded by color degradation and marine snow artifacts. Generating underwater images from atmospheric RGB-D images based on physical models provides pairs of real-world ground-truth and degraded images. Our synthetic approach generates a large quantity of the pairs, enabling effective training of deep neural networks and objective image quality assessment. Through benchmark experiment with some datasets and image enhance methods, we validate that our dataset can improve the image enhancement performance. Our dataset, which is publicly available, contributes to the development in underwater image processing.
@@ I Introduction
Underwater image enhancement is one of the extreme image restoration challenges. The primary hurdle lies in the absence of ground-truth (atmospheric) images. Unlike image processing in atmospheric conditions, obtaining clean atmospheric counterparts for underwater images is not feasible. This absence of ground truth makes it difficult to objectively assess restoration quality. Furthermore, the lack of pairs of ground-truth and degraded images hampers neural network training, as traditional loss functions (e.g., peak signal to noise ratio (PSNR) and structual similarity (SSIM)) evaluate the error between these pairs [23].
Various datasets and enhancement methods have tackled to address this lack of ground-truth images[24, 1, 13, 25, 14, 26, 27, 28]. However, most datasets are not publicly available and are often oversimplified color degradation processes (details are described in Section II-A). Moreover, they overlook a critical aspect of degradation: Marine snow. Marine snow, visible as white dots in digital images as shown in Fig. 1, is caused by light reflecting off particles like organic matter and sand. This phenomenon degrades overall image quality, yet its removal techniques are still nascent, often limited to classical median-filtering approaches [29, 30, 31]. The lack of datasets has also made it difficult to develop effective removal techniques for marine snow.
@@
In this paper, we propose a new dataset, the physics-inspired synthesized underwater image dataset (PHISWID), providing pairs of ground-truth images, and underwater images with color degradation and marine snow artifacts [32, 33]. Our dataset is based on physical observation models which enable to synthesis underwater image from real-world atmospheric images. For the physical observation models, the light absorption and scattering model in underwater is used to model color shift [34, 35] and the light scattering model for underwater particle to marine snow [36]. We apply these models to an atmospheric RGB-D image to generate its synthetic underwater image, obtaining a pair of the ground-truth and underwater images.
@@
Our dataset is validated with various image enhancement methods and some existing datasets. We show that the learning-based methods trained with our dataset outperforms the ones trained with the other datasets [11, 37], even with basic U-Net [38]. Our benchmark results suggest a potential contribution of our dataset, which is made publicly available, to development of the underwater image processing [^1] https://github.com/reina0112/PHISMID-PHISWID..
@@
PHISWID (proposed) Color shift & Marine snow 2264 Physics-based synthesis
@@ II-B Marine Snow Removal
  • Ground truth clean atmospheric images, i.e., target images.
@@ III-A1 Algorithm Overview
8:        for $k=1$ to $10$ do
9:           Set $d_{\text{vert}}$ within $0.5\leq d_{\text{vert}}\leq d_{\text{vert},\text{max}}$ randomly
10:           Set $B_{c}$ within $B_{c,\text{min}}\leq B_{c}\leq B_{c,\text{max}}$ randomly
11:           Randomly select a camera $r$
12:           $\mathrm{SR}(\lambda)$leftarrow $D_{\mathrm{SR}(\lambda),r}$
13:           Calculate equation (1)–(5) for all pixels and get $U_{c}$
14:           Put $U_{c}$ into $D_{\mathcal{W}_{t},\text{synthesized}}$
@@
Our color shift model considers attenuation of light due to underwater condition and spectral difference of camera property. For modeling the attenuation, we use a model proposed by Jerlov [34] which presents attenuation of light at each wavelength (400 to 700 nm). The attenuation coefficients are varied for water types, which Jerlov classified into ten as shown in Table II. In addition to the attenuation property, the color also depends on the spectrum response which is an inherent property of a camera and lens. Therefore, our color shift synthesis takes into account the attenuation and spectral response to simulate the color shift in underwater environment.
@@
As a result, our algorithm creates ten sets of synthetic images, each corresponds to one water type.
@@ III-B Synthesizing Color Shift
The attenuation coefficients for ten water types $D_{w,\beta(\lambda)}$ (see Algorithm 1), are given by [53]. We also use the data of solar irradiance on the water surface [54] as $\Theta_{0}(\lambda)$. The camera spectrum responses are also given from the dataset of 28 digital cameras [55]. The other parameters are set to $d_{\text{vert},\text{max}}=1$, $\{B_{R,\text{min}},B_{G,\text{min}},B_{B,\text{min}}\}=\{0.4,0.7,0.7\}$, and $\{B_{R,\text{max}},B_{G,\text{max}},B_{B,\text{max}}\}=\{0.5,0.8,0.8\}$ [1]. We randomly determine five sets of $B_{c}$ and $d_{\text{vert}}$ from the uniform distribution in the range of $B_{c,\text{min}}\leq B_{c}\leq B_{c,\text{max}}$ and $0.5\leq d_{\text{vert}}\leq d_{\text{vert},\text{max}}$, respectively. As a result, 50 underwater images are generated from a single atmospheric image with the ten water types and the five sets of $B_{c}$ and $d_{\text{vert}}$.
@@ III-C2 Modeling Marine Snow Artifacts
    Determining coordinates of particles: As shown in Fig. 5, suppose that the camera is positioned parallel to the $(x,y)$-plane and at a sufficiently large distance from the scene located in the three-dimensional Euclidean space. The three-dimensional coordinates of the particles are uniform-randomly set. The number of marine snow particles is given for each image.
@@ IV PHISWID Specifications
Our developed dataset, PHISWID, contains $2264$ image pairs, all having a pixel resolution of $384\times 384$. All original atmospheric RGB-D images used are collected from NYD-RGB dataset [32] and an outdoor image dataset [33]. An image pair contains one original atmospheric image and one synthesized underwater image degraded by color shift and marine snow artifacts. We randomly set water type for each image. Image examples in PHISWID are shown in Fig. 7e.
@@
Most parameters are determined according to our preliminary observations of real underwater images: They are summarized in Appendix B. Further examples of PHISWID images are provided in Appendix C.
@@ V-A Setup
U-Net(P) U-Net PHISWID
U-Net(P’) U-Net PHISWID w/o marine snow
Deep WaveNet(P) Deep WaveNet PHISWID
U-Net(U) U-Net UIEB
Deep WaveNet(U) [43] Deep WaveNet UIEB
WaterNet [11] WaterNet UIEB
U-Net(L) U-Net LSUI
@@
We divided PHISWID to 2184 training image pairs and 80 test image pairs. Underwater image enhancement performance is compared with methods summarized in Table IV. The parameters of the alternative methods are basically followed to the original researches. We use WaterNet [11], Deep WaveNet [43], and HLRP [28], which are widely used for underwater image enhancement and restoration. We also trained Deep WaveNet with the proposed PHISWID to compare dataset qualities.
Furthermore, we train basic U-Nets with PHISWID, UIEB [11], and LSUI [37]. The structure of the U-Net is shown in Appendix D. For a comparison purpose, we also train the same U-Net with the dataset without marine snow artifacts, i.e., the dataset only applied our color shift model.
@@ V-B Results
The experimental results are shown in Fig. 10 and the objective performances are summarized in Table V. Both PSNRs and SSIMs of the neural networks trained with PHISWID outperform those with the existing datasets. Fig. 10 also visualizes such differences: The methods trained with PHISWID restore images like their atmospheric counterparts. On the other hand, marine snow artifacts remain and blueish images are still generated by methods trained with the alternative datasets. U-Net(P’) can remove bluish color shift, however, it cannot remove marine snow artifacts.
@@
Method PSNR SSIM
U-Net(P) 23.97 0.714
U-Net(P’) 21.68 0.654
Deep WaveNet(P) 28.09 0.818
U-Net(U) 20.05 0.162
Deep WaveNet(U) 21.24 0.166
Water Net 20.89 0.439
U-Net(L) 21.33 0.319
U-shape 24.49 0.413
HLRP 15.59 0.118
Synthesized image 19.51 -0.010
(a) Test images
in PHISWID
(b) U-Net(P)
(c) U-Net(P’)
(d) Deep
WaveNet(P)
(e) U-Net(U)
(f) Deep
WaveNet(U) [43]
(g) WaterNet [11]
(h) U-Net(L)
(i) U-shape [37]
(j) HLRP [28]
Figure 10: PHISWID results.
@@ VI-A Real-world Underwater Image Enhancement
Fig. 11 shows examples of underwater image enhancement for real-world images. We compare the methods used in Section V-A. As visual evaluation, methods trained with our proposed dataset remove color shift sufficiently. Especially, a significant difference between PHISWID and the other datasets can be observed in background regions. The PHISWID-based methods can remove the bluish color shift in the background while the others fails.
According to the removal of marine snow artifacts, the PHISWID-based methods succeeded remove them. U-Net(P’) and the methods trained with the existing datasets without marine snow artifacts naturally fail to remove the artifacts. This suggests that PHISWID is beneficial for real-world underwater image enhancement.
(a) Original
underwater images
(b) U-Net(P)
(c) U-Net(P’)
(d) Deep
WaveNet(P)
(e) U-Net(U)
(f) Deep
WaveNet(U)
(g) Water Net
(h) U-Net(L)
(i) U-shape
(j) HLRP
@@ VI-B Limitations
Although our PHISWID are carefully designed based on physical measurement processes, we still assume many parameters and settings: This is one of the limitations.
@@
With regard to PHISWID, if the light scattering is very different from the training images, color shift, particularly in the distant backgrounds, could not be well restored as shown in Fig. 12. Its possible reason is that the RGB-D images used in PHISWID are recorded in indoor situations which results in limited horizontal distances (i.e., depth) in scenes. Therefore it may not generate color shifts of distant backgrounds.
Our future work includes to update PHISWID for covering wider degradation patterns like hazes.
Figure 12: Limitations. Top: Original images. Bottom: Restoration results by U-Net trained with PHISWID.

### v3

# PHISWID: Physics-Inspired Underwater Image Dataset Synthesized from RGB-D Images

@@ Abstract
This paper introduces the physics-inspired synthesized underwater image dataset (PHISWID), a dataset tailored for enhancing underwater image processing through physics-inspired image synthesis. For underwater image enhancement, data-driven approaches (e.g., deep neural networks) typically demand extensive datasets, yet acquiring paired clean atmospheric images and degraded underwater images poses significant challenges. Existing datasets have limited contributions to image enhancement due to lack of physics models, publicity, and ground-truth atmospheric images. PHISWID addresses these issues by offering a set of paired atmospheric and underwater images. Specifically, underwater images are synthetically degraded by color degradation and marine snow artifacts from atmospheric RGB-D images. It is enabled based on a physics-based underwater image observation model. Our synthetic approach generates a large quantity of the pairs, enabling effective training of deep neural networks and objective image quality assessment. Through benchmark experiments with some datasets and image enhancement methods, we validate that our dataset can improve the image enhancement performance. Our dataset, which is publicly available, contributes to the development in underwater image processing.
@@ I Introduction
Underwater image enhancement is one of the extreme image restoration challenges. In underwater, light (especially their red components) is strongly absorbed and scattered, rather than that in atmospheric scenes. This is due to two “depths” in underwater scenes: 1) horizontal distances between the camera and scene (this is the typical “depth” in regular image processing) and 2) water depth. Both of them affect light absorption and scattering properties, which result in shifting the color of images in underwater [23]. Images are also degraded due to marine snow [24], whose artifacts are visible as white dots due to light reflecting off particles like organic matter and sand. Its example is shown in Fig. 1. As a result, underwater image degradation is significantly different from atmospheric images, which results in requirements for image enhancement methods designed for underwater scene.
Many underwater image enhancement methods, both model- and deep learning-based ones, have been proposed [25, 10, 26, 27, 28, 29, 30, 31]. While their building blocks vary, most works aim to the same goal—producing clean underwater images. They may be enough for entertainment purposes, however, in many situations like rescue, defense, bioprospecting/resource exploration, we often need enhanced underwater images looks like atmospheric images. For example, we can use an object detection system designed for regular atmospheric images as-is when underwater images are restored like atmospheric ones. Fig. 2 shows an example of object detection for a real underwater image and enhanced ones. We simply use a YOLOv5 pretrained model [32] for the object detection experiment. As can be seen, a person cannot be detected with the original underwater image. All of the enhanced images help to detect the person, but the confidence scores are different: Reducing blueish color shift will contribute the accurate detection.
@@
The primary hurdle for underwater image enhancement, especially for yielding atmospheric-like images, lies in the absence of ground-truth images. Unlike regular image processing, obtaining clean atmospheric counterparts for underwater images is not feasible. The lack of pairs of ground-truth and degraded images hampers neural network training, as traditional loss functions (e.g., peak signal to noise ratio (PSNR) and structual similarity (SSIM)) evaluate the error between these pairs [33].
Various underwater image datasets and enhancement methods have tackled to address this lack of ground-truth images [34, 11, 13, 35, 14, 36, 37, 38]. However, we face the following three problems:
  1. i.
    Many datasets are not publicly available [35, 12, 39].
  2. ii.
    They often produce clean images as reverse engineering: Ground-truth images are obtained by (manual or automatic) enhancement of underwater images. This process does not take into account the above-mentioned two “depths” information in underwater since it is not accompanied with the images. Moreover, enhanced underwater images, i.e., cleaned underwater images, are used as ground-truth [10, 26, 40].
  3. iii.
    Many methods overlook a critical aspect of marine snow. This phenomenon degrades overall image quality, yet its removal techniques are often limited to classical median-filtering approaches [41, 42, 43].
@@
To tackle the above-mentioned challenges, this paper propose a new dataset, the physics-inspired synthesized underwater image dataset (PHISWID), providing pairs of real atmospheric ground-truth images and synthesized underwater images with color degradation and marine snow artifacts taking into account the two-depths information [44, 45]. Our dataset is based on physical observation models which enable to synthesis underwater image from real-world RGB-D images. For the physical observation models, the light absorption and scattering model in underwater is used to model color shift [23, 46] and the light scattering model for underwater particle to marine snow [47]. We apply these models to an atmospheric RGB-D image to generate its synthetic underwater image, obtaining a pair of the ground-truth and underwater images.
@@
Our dataset is validated with various image enhancement methods and some existing datasets. We show that the learning-based methods trained with our dataset outperforms the ones trained with the other datasets [10, 26], even with basic Transformer [48]. Our benchmark results suggest a potential contribution of our dataset, which is made publicly available, to development of the underwater image processing [^1] https://github.com/reina0112/PHISMID-PHISWID..
@@
(a) Original image
(b) Enhanced result of PHISWID.
(c) Enhanced result of LSUI.
(d) Enhanced result of UIEB.
Figure 2: Object detection results of a test image in the EUVP dataset [31]. We use a YOLOv5 pretrained model [32] for the object detection experiment. The blue squares are estimated bounding boxes and the numbers are confidence score ($[0,1]$). (a) original image (an object is not detected). (b) enhanced image based on PHISWID. (c) enhanced image based on LSUI. (d) enhanced image based on UIEB. The enhancement method is the same as that in the experiment in Section V.
@@
PHISWID (proposed) Color shift & Marine snow 4195 Physics-based synthesis
@@
The preliminary version of this paper was presented in [11]. This paper extends it by 1) incorporating marine snow artifacts with physics-based observation models in the dataset, 2) improving the image synthesis process and enhancing resolutions, and 3) providing comprehensive comparisons with alternative datasets and restoration methods.
@@ II-B Marine Snow Removal
  • Ground-truth clean atmospheric images, i.e., target images.
@@ III-A1 Algorithm Overview
8:        Set $d_{\text{vert}}$ within $0.5\leq d_{\text{vert}}\leq d_{\text{vert},\text{max}}$ randomly
9:        Set $B_{c}$ within $B_{c,\text{min}}\leq B_{c}\leq B_{c,\text{max}}$ randomly
10:        Randomly select a camera $r$
11:        $\mathrm{SR}(\lambda)$leftarrow $D_{\mathrm{SR}(\lambda),r}$
12:        Calculate equation (1)–(5) for all pixels and get $U_{c}$
13:        Put $U_{c}$ into $D_{\mathcal{W}_{t},\text{synthesized}}$
14:     until Generating $U_{c}$ from all images in $D_{\text{RGB-D}}$
@@
Our color shift model considers attenuation of light due to underwater condition and spectral difference of camera property. For modeling the attenuation, we use a model proposed by Jerlov [23] which presents attenuation of light at each wavelength (400 to 700 nm). The attenuation coefficients are varied for water types, which Jerlov classified into ten as shown in Table II. We omit three coastal water types, 5C, 7C, and 9C, from our dataset because objects are almost invisible for these types. In addition to the attenuation property, the color also depends on the spectrum response which is an inherent property of a camera and lens. Therefore, our color shift synthesis takes into account the attenuation and spectral response to simulate the color shift in underwater environment.
@@
Our algorithm creates seven sets of synthetic images, each corresponds to one water type.
@@ III-B Synthesizing Color Shift
The attenuation coefficients for seven water types $D_{w,\beta(\lambda)}$ (see Algorithm 1), are given by [54]. We also use the data of solar irradiance on the water surface [55] as $\Theta_{0}(\lambda)$. The camera spectrum responses are also given from the dataset of 28 digital cameras [56]. The other parameters are set to $d_{\text{vert},\text{max}}=1$, $\{B_{R,\text{min}},B_{G,\text{min}},B_{B,\text{min}}\}=\{0.4,0.7,0.7\}$, and $\{B_{R,\text{max}},B_{G,\text{max}},B_{B,\text{max}}\}=\{0.5,0.8,0.8\}$ [11]. We randomly determine one of $B_{c}$ and $d_{\text{vert}}$ from the uniform distribution in the range of $B_{c,\text{min}}\leq B_{c}\leq B_{c,\text{max}}$ and $0.5\leq d_{\text{vert}}\leq d_{\text{vert},\text{max}}$, respectively. As a result, 7 underwater images are generated from a single atmospheric image with the seven water types and the one set of $B_{c}$ and $d_{\text{vert}}$.
@@ III-C2 Modeling Marine Snow Artifacts
    Particle Setting: As shown in Fig. 6, suppose that the camera is positioned parallel to the $(x,y)$-plane and at a sufficiently large distance from the scene located in the three-dimensional Euclidean space. The three-dimensional coordinates of the particles are uniform-randomly set. The number of marine snow particles is given for each image.
@@ IV PHISWID Specifications
Our developed dataset, PHISWID, contains $4195$ image pairs, all having a pixel resolution of $1344\times 756$. All original atmospheric RGB-D images used are collected from the large-scale RGB-D database [59, 60, 61, 62, 63]. An image pair contains one original atmospheric image and one synthesized underwater image degraded by color shift and marine snow artifacts. We randomly set water type for each image. Image examples in PHISWID are shown in Fig. 8t.
@@
Most parameters are determined according to our preliminary observations of real underwater images: They are summarized in Appendix B.
Figure 9: *
@@
(a) Test images
(b) Ground-truth images
(c) Trans(P)
(d) Trans(P’)
(e) DWN(P)
(f) Trans(U)
(g) DWN(U)
(h) WN
(i) Trans(L)
(j) U-shape
(k) HLRP
Figure 11: Image enhancement results with PHISWID test images. From up to down, we use different water types
@@ V-A Setup
Trans(P) Transformer PHISWID
Trans(P’) Transformer PHISWID w/o marine snow
DWN(P) Deep WaveNet PHISWID
Trans(U) Transformer UIEB
DWN(U) [25] Deep WaveNet UIEB
WN [10] WaterNet UIEB
Trans(L) Transformer LSUI
@@
We divided PHISWID to 4125 training image pairs and 70 test image pairs. Underwater image enhancement performance is compared with various methods summarized in Table IV. The parameters of the alternative methods are basically followed to the original researches. We use WaterNet (abbreviated as WN) [10], Deep WaveNet (abbreviated as DWN) [25], U-shepe Transformer (abbreviated as U-shape) [26] and HLRP [38], which are widely used for underwater image enhancement and restoration. WN, DWN(U), and U-shape are trained with their own datasets.
To compare dataset qualities, we trained DWN with the proposed PHISWID. Furthermore, we train Transformers with PHISWID (abbreviated as Trans(P)), UIEB (abbreviated as Trans(U)) [10], and LSUI (abbreviated as Trans(L)) [26]. The structure of the Transformer is shown in Appendix C. For a comparison purpose, we also train the same Transformer with the dataset without marine snow artifacts, i.e., the dataset only applied our color shift model (abbreviated as Trans(P’)).
For objective quality comparison, we compute the average PSNRs and SSIMs over the test datasets. We also show naturalness image quality evaluator (NIQE) [64] as a non-reference quality metric. NIQE fits our purpose of the paper because it returns a low score when images are natural taken in atmospheric scenes.
@@ V-B Results
The experimental results are shown in Fig. 11 and the objective performances are summarized in Table V. All of the metrics (PSNRs, SSIMs, and NIQEs) of the neural networks trained with PHISWID outperform those with the existing datasets. Fig. 11 visualizes such differences: The methods trained with PHISWID restore images like their atmospheric counterparts. On the other hand, marine snow artifacts remain and blueish images are still generated by methods trained with the alternative datasets. Trans(P’) can remove bluish color shift, however, it cannot remove marine snow artifacts.
@@
[…]
Trans(P) 21.00 0.793 2.86
Trans(P’) 19.10 0.748 3.69
DWN(P) 21.10 0.671 3.11
Trans(U) 14.69 0.286 4.13
DWN(U) 16.60 0.359 3.67
WN 17.06 0.359 3.76
Trans(L) 15.06 0.363 4.05
U-shape 14.88 0.368 3.11
HLRP 7.10 0.062 NaN
Synthesized image 13.67 0.120 4.02
(a) Real images
(b) Trans(P)
(c) Trans(P’)
(d) DWN(P)
(e) Trans(U)
(f) DWN(U)
(g) WN
(h) Trans(L)
(i) U-shape
(j) HLRP
(a) Real images
(b) Trans(P)
(c) Trans(P’)
(d) DWN(P)
(e) Trans(U)
(f) DWN(U)
(g) WN
(h) Trans(L)
(i) U-shape
(j) HLRP
Figure 13: Underwater image enhancement results for real-world images in UIEB challenges.
(a) Real images
(b) Trans(P)
(c) Trans(P’)
(d) DWN(P)
(e) Trans(U)
(f) DWN(U)
(g) WN
(h) Trans(L)
(i) U-shape
(j) HLRP
Figure 14: Enlarged portion of marine snow removal.
TABLE VI: Average NIQE over real images.
[…]
Trans(P) 4.07
Trans(P’) 4.01
DWN(P) 3.88
Trans(U) 4.30
DWN(U) 4.21
WN 4.64
Trans(L) 4.57
U-shape 4.97
HLRP NaN
Original image 4.62
@@ VI-A Real-world Underwater Image Enhancement
We compare the methods used in Section V-A through real-world underwater images.
Fig. 12 shows underwater image enhancement results for real-world images. All original images are collected from flickr [^4] https://www.flickr.com/ under a Creative Commons AttributionNonCommercial-ShareAlike 2.0 Generic (CC BY-NC-SA 2.0) License. We also show results for image taken from the UIEB challenging set [10] in Fig. 13. We calculate the average NIQE of 66 real-world images, including Fig. 12 and UIEB challenging sets, is summarized in Table VI.
As shown in the figures, methods trained with PHISWID remove color shift sufficiently. A significant difference between PHISWID and the other datasets can be observed in background regions. The PHISWID-based methods have a strong effect to remove the bluish color shift in the background.
The PHISWID-based methods also remove marine snow artifacts and small bubbles similar to marine snow artifacts. Fig. 14 shows enlarged portions of the enhancement results. In contrast, Trans(P’) and the methods trained with the existing datasets fail to remove the artifacts.
@@ VI-B Limitations
Although our PHISWID are carefully designed based on physical measurement processes, we still assume many parameters and settings as shown in Table VII: This is one of the limitations.
@@
Depths beyond our assuming ranges also affect the overall qualities. If the light scattering is very different from the training images, color shift, particularly in the distant backgrounds, could not be well restored as shown in Fig. 15. A possible reason for this low performance is that the RGB-D images used in PHISWID are recorded in indoor situations which results in limited horizontal distances (i.e., depth) in scenes. Therefore it may not appropriately restore color shifts of distant backgrounds.
Our future work includes to update PHISWID for covering broader distance and degradation patterns.
Figure 15: Limitations. Top: Original images. Bottom: Restoration results by Trans(P).